You are in line at a ticket window. There are two more people ahead of you than there are behind you. In the entire line, there are three times as many people as there are behind you. How many people are ahead of you in line?
Answer by xcentaur(357)


You are in line at a ticket window.
Cool,then?



There are two more people ahead of you than there are behind you.

Let the people behind you be x number of people.
Then the number of people ahead of you are (x+2).



In the entire line, there are three times as many people as there are behind you.

so the number of people behind you is x
the entire line has three times x people = 3x


How many people are ahead of you in line?
Look at it this way:
you have (x+2) people ahead of you
then you are standing there
then you have x number of people behind you
**so all these people put together make the total number of people in the line.
total number of people=(x+2)+1+x=(2x+3)
but the question tell us that total number of people in the line is 3x
**so we get,
3x=2x+3
3x-2x=3
x=3


remember x stands for the people behind you in the line
so you have 3 people behind you in the line
the question says there are (x+2) people ahead of you
so there are (3+2)=5 people ahead of you in the line.


Hope this helps,
Prabhat